OpenStudy (anonymous):

would the answer of (a − b)(a − b) be a^2 - ab^2 - b^2

OpenStudy (anonymous):

FOIL First: a*a = a^2 Outside = -ab Inside = -ab Last = -b(-b) = +b^2 What do we do with the -ab and the -ab?

OpenStudy (anonymous):

Tkink -x + -x = -2x

OpenStudy (anonymous):

I know to combine itt

OpenStudy (anonymous):

@chungerforever remember when we had \(5x+5x=10x\)? don't change the exponents here either

OpenStudy (anonymous):

\(-ab-ab=?\)

OpenStudy (anonymous):

oh minus! so just ab?

OpenStudy (anonymous):

\(-10-10=?\)

OpenStudy (anonymous):

oh I understand, so the answer is just a^2 - b^2?

OpenStudy (anonymous):

hold the phone, careful arithmetic \(-10-10=?\)

OpenStudy (anonymous):

hint, it is not 0

OpenStudy (anonymous):

-10?

OpenStudy (anonymous):

don't fret, we will work it out, but i was asking an arithmetic question \(-10-10=-20\)

OpenStudy (anonymous):

\(-100-100=?\)

OpenStudy (anonymous):

-200?

OpenStudy (anonymous):

yes that is it !

OpenStudy (anonymous):

\(-7-7=?\)

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

ok good now we can move on to variables

OpenStudy (anonymous):

\[-10-10=-20\] \[-100-100=-200\] \[-7-7=-14\] \[-12-12=-24\] and so on how about \[-x-x\]? well this must be the same as the others, namely \[-x-x=-2x\]

OpenStudy (anonymous):

so -2ab?

OpenStudy (anonymous):

notice that the exponent does not change spend $20, spend another $20 you have spent \(2\times 20=40\) yes you have it !

OpenStudy (anonymous):

so "final answer" for \((a-b)(a-b)\) is ??

OpenStudy (anonymous):

a^2 - 2ab - b^2?

OpenStudy (anonymous):

yes that is it!

OpenStudy (anonymous):

thank you so much! and sorry I was taking a while to reply. But thanks again I appreciate it (:

OpenStudy (anonymous):

yw good work, hope you are understanding a bit more about combining like terms